OpenStudy (anonymous):

Explain why (–4x)^0 = 1, but –4x^0 = –4.

OpenStudy (lgbasallote):

becaus x is the one raised to zero

OpenStudy (angela210793):

cause (-4x)^0=-4^0*x^0=1*1=1 -4x^0 means tht only x is in the 0th power---> -4*1=-4

OpenStudy (amistre64):

notation matters

OpenStudy (anonymous):

x^0 means 1 where x refers any variable even a thing or a name or anything else! hence replace x with(-4x) hence (–4x)^0 = 1